OpenStudy (anonymous):

6. The length of a rectangle is 3 cm more than twice its width. If the perimeter of the rectangle is 54 cm, find the length of the rectangle.

OpenStudy (anonymous):

let the width be 'x' and length be '2x+3' 2(l+b) = 54 2x+3+x = 54/2 3x = 27-3 x = 24/3 x = 8 so,width=8 length = 2(8)+3 = 16+3 = 19